# compat=1.0.0

### v1

@@ 3 Numerical results and updated constraints on Higgs boson self-coupling
$\delta\sigma^{\kappa_{\lambda}}_{\rm ggF,EW}$ $=(0.075\kappa_{\lambda_{3}}^{4}-0.158\kappa_{\lambda_{3}}^{3}-0.006\kappa_{\lambda_{3}}^{2}\kappa_{\lambda_{4}}-0.058\kappa_{\lambda_{3}}^{2}+0.070\kappa_{\lambda_{3}}\kappa_{\lambda_{4}}-0.149\kappa_{\lambda_{4}})~{}~{}{\rm fb}$
@@
$\delta\sigma_{\rm VBF,EW}^{\kappa_{\lambda}}$ $=(0.0215\kappa_{\lambda_{3}}^{4}-0.0324\kappa_{\lambda_{3}}^{3}-0.0019\kappa_{\lambda_{3}}^{2}\kappa_{\lambda_{4}}-0.0043\kappa_{\lambda_{3}}^{2}$
$~{}~{}~{}+0.0151\kappa_{\lambda_{3}}\kappa_{\lambda_{4}}-0.0211\kappa_{\lambda_{4}})~{}~{}{\rm fb}$ (12)
for the VBF channel. We have computed all the $\mathcal{O}(\lambda_{\rm 3H}^{i}),i\geq 2$ contributions in the amplitude. The above $\kappa_{\lambda_{3}}^{2}$ terms arise because we want to keep the cancellation relation between the $\mathcal{O}(\lambda_{\rm 3H})$ and $\mathcal{O}(1)$ amplitudes at LO. The cubic $\kappa_{\lambda_{3}}^{3}$ and quartic $\kappa_{\lambda_{3}}^{4}$ terms appear for the first time up to this perturbative order. Though their coefficients are rather small, they provide notable corrections to the cross section if $\kappa_{\lambda_{3}}$ is chosen much larger than 1. As seen from table 1, the $\lambda$ dependent corrections in the ggF (VBF) channel reach $91\%$ ($82\%$) of the LO cross section for $\kappa_{\lambda_{3}}=6$.
$\kappa_{\lambda_{3}}$ $\kappa_{\lambda_{4}}$ ggF VBF
@@
Table 1: Cross sections (in fb) of ggF and VBF Higgs boson pair production for different $\kappa_{\lambda_{3}}$ and $\kappa_{\lambda_{4}}$ at the 13 TeV LHC.
@@
In addition, there is a new dependence on the quartic Higgs self-coupling $\lambda_{\rm 4H}$. Because this dependence is only linear and the corresponding coefficients are small, their contributions are negligible. From table 1, it can be observed that the cross section varies by $0.6\%$ when $\kappa_{\lambda_{4}}$ changes from 1 to 6 while keeping $\kappa_{\lambda_{3}}=6$. As a consequence, we do not expect that a meaningful constraint on the quartic Higgs self-coupling can be extracted from Higgs pair production at the 13 TeV LHC.
@@
We can make a comparison with the results in Refs. [65, 31]. The authors of these papers have obtained similar expressions for the cross sections in the ggF channel. However, they have assumed that the triple and quartic Higgs self-couplings are modified by one dimension-six and one dimension-eight operators. They performed calculations, especially renormalization, in terms of the coefficients of higher-dimensional operators, and then transformed the results onto the basis of $\kappa_{\lambda_{3}}$ and $\kappa_{\lambda_{4}}$. These results can not be directly compared with the experimental analysis in the $\kappa$ framework.
@@
Figure 3: Cross sections of Higgs boson pair production as a function of $\kappa_{\lambda_{3}}=\kappa_{\lambda_{4}}=\kappa_{\lambda}$ at the 13 TeV LHC including both ggF and VBF processes. The black line represents the LO result, and the green line denotes the result with (N)NNLO QCD corrections in the ggF (VBF) channel. The red line indicates the result with higher power dependence on the Higgs boson self-coupling.
@@
In figure 3, we show different perturbative predictions for the cross section of Higgs boson pair production at the 13 TeV LHC as a function of $\kappa_{\lambda_{3}}=\kappa_{\lambda_{4}}=\kappa_{\lambda}$. It is obvious that higher-order perturbative corrections dramatically change the function form. The current experimental upper limit by the ATLAS (CMS) collaborations on $\kappa_{\lambda}$ is $6.6$ (6.49) based on the theoretical calculations at QCD NNLO in the ggF channel and NNNLO in the VBF channel; see table 1. Taking $\delta\sigma^{\kappa_{\lambda}}_{\rm EW}$ corrections into account and assuming that the QCD and EW corrections are factorizable, the upper limit is narrowed down to 5.4 (5.37). These limits are almost the same when keeping $\kappa_{\lambda_{4}}=1$. If the scale uncertainties are considered [66], the upper limit spans in the range $(6.5,6.8)$ in the ATLAS result, which would decrease to $(5.4,5.6)$ after including higher power dependence. In the CMS result, the upper limit changes from (6.40, 6.67) to (5.31, 5.48). The lower limits are only modified slightly.
@@
Lastly, the Higgs boson pair invariant mass distributions in the ggF channel are shown in figure 4. The peak position moves from 400 GeV to 260 GeV when $\kappa_{\lambda}$ changes from 1 to 6, which indicates the Higgs bosons tend to be produced with very small velocity in the case of large $\kappa_{\lambda}$ values. For $\kappa_{\lambda}=3$, there is a bump around $m_{HH}=450$ GeV. These features may help to set optimal cuts in the experimental analysis. From this figure, we also observe that the $\lambda$ dependent corrections have a great impact on the distributions, especially in the small $m_{HH}$ region, and therefore they should be included in future studies.
@@
Figure 4: The distributions of Higgs boson pair invariant mass at LO and with $\delta\sigma_{\rm EW}^{\kappa_{\lambda}}$ corrections at the 13 TeV LHC. We have used $\kappa_{\lambda_{3}}=\kappa_{\lambda_{4}}=\kappa_{\lambda}$.
@@ 4 Conclusion
The exact shape of the Higgs potential is one of the most important mysteries in particle physics. The current limits on the Higgs self-coupling are predominantly derived based on the assumption that the cross section of the Higgs boson pair production is a quadratic function of the self-coupling. We find that the function form should be generalized to include quartic and cubic dependence on the self-coupling which arise due to higher-order corrections induced by virtual Higgs bosons.
We adopt a proper renormalization procedure to keep track of all the contributions from Higgs self-couplings. The complicated two-loop squared amplitudes are computed numerically and a grid is generated for the calculation of the cross section. We present numerical contributions of the quartic and cubic terms of the trilinear Higgs self-coupling in the cross section at the LHC. With this refined functional form, we demonstrate that the upper limit on the trilinear Higgs self-coupling normalized by the SM value is reduced from 6.6 (by ATLAS) and 6.49 (by CMS) to 5.4 and 5.37, respectively. This improvement is achieved without more data being analyzed, highlighting the importance of including higher power dependence on the Higgs self-coupling in the cross section.
Moreover, we find it hard to obtain any useful constraint on the quartic Higgs self-coupling from Higgs boson pair production. One may resort to triple Higgs boson production. This requires a collider with higher energies than the LHC. In the future, it would also be interesting to investigate the impact of even higher power dependence to see if the limit could be strengthened further.
@@ Acknowledgments
We would like to thank Huan-Yu Bi, Yan-Qing Ma, and Huai-Min Yu for comparing their numerical results of the two-loop amplitudes. We also thank Shan Jin, Yefan Wang and Lei Zhang for helpful discussion. This work was supported in part by the National Science Foundation of China under grant No. 12275156, No. 12321005, No. 12375076 and the Taishan Scholar Foundation of Shandong province (tsqn201909011).

### v2

@@ 3 Numerical results and improved constraints on the Higgs boson self-coupling
$\delta\sigma^{\kappa_{\lambda}}_{\rm ggF,EW}$ $=(0.075\kappa_{\lambda_{\rm 3H}}^{4}-0.158\kappa_{\lambda_{\rm 3H}}^{3}-0.006\kappa_{\lambda_{\rm 3H}}^{2}\kappa_{\lambda_{\rm 4H}}-0.058\kappa_{\lambda_{\rm 3H}}^{2}$
$~{}~{}~{}+0.070\kappa_{\lambda_{\rm 3H}}\kappa_{\lambda_{\rm 4H}}-0.149\kappa_{\lambda_{\rm 4H}})~{}~{}{\rm fb}$ (12)
@@
$\delta\sigma_{\rm VBF,EW}^{\kappa_{\lambda}}$ $=(0.0215\kappa_{\lambda_{\rm 3H}}^{4}-0.0324\kappa_{\lambda_{\rm 3H}}^{3}-0.0019\kappa_{\lambda_{\rm 3H}}^{2}\kappa_{\lambda_{\rm 4H}}-0.0043\kappa_{\lambda_{\rm 3H}}^{2}$
$~{}~{}~{}+0.0151\kappa_{\lambda_{\rm 3H}}\kappa_{\lambda_{\rm 4H}}-0.0211\kappa_{\lambda_{\rm 4H}})~{}~{}{\rm fb}$ (13)
for the VBF channel. We have computed all the $\mathcal{O}(\lambda_{\rm 3H}^{i}),i\geq 2$ contributions in the amplitude. The above $\kappa_{\lambda_{\rm 3H}}^{2}$ terms arise because we want to keep the cancellation relation between the $\mathcal{O}(\lambda_{\rm 3H})$ and $\mathcal{O}(1)$ amplitudes at LO. The cubic $\kappa_{\lambda_{\rm 3H}}^{3}$ and quartic $\kappa_{\lambda_{\rm 3H}}^{4}$ terms appear for the first time up to this perturbative order. Though their coefficients are rather small, they provide notable corrections to the cross section if $\kappa_{\lambda_{\rm 3H}}$ is chosen much larger than 1. As seen from Table 1, the $\lambda$ dependent corrections in the ggF (VBF) channel reach $91\%$ ($82\%$) of the LO cross section for $\kappa_{\lambda_{\rm 3H}}=6$.
$\kappa_{\lambda_{\rm 3H}}$ $\kappa_{\lambda_{\rm 4H}}$ ggF VBF
@@
Table 1: Cross sections (in fb) of ggF and VBF Higgs boson pair production for different values of $\kappa_{\lambda_{\rm 3H}}$ and $\kappa_{\lambda_{\rm 4H}}$ at the 13 TeV LHC.
@@
In addition, there is a new dependence on the quartic Higgs self-coupling $\lambda_{\rm 4H}$. Because this dependence is only linear and the corresponding coefficients are small, their contributions are negligible. From Table 1, it can be observed that the cross section varies by $0.6\%$ when $\kappa_{\lambda_{\rm 4H}}$ changes from 1 to 6 while keeping $\kappa_{\lambda_{\rm 3H}}=6$. As a consequence, we do not expect that a meaningful constraint on the quartic Higgs self-coupling can be extracted from Higgs boson pair production at the LHC.
@@
We can make a comparison with the results in Refs. [71, 31]. The authors of these papers have obtained similar expressions for the cross sections in the ggF channel. However, they have assumed that the triple and quartic Higgs self-couplings are modified by one dimension-six and one dimension-eight operators  [^3] There are new physics scenarios, e.g, the Nambu-Goldstone Higgs potential [72], in which the modification of the Higgs potential can not be described by higher-dimensional operators but can be accommodated in the $\kappa$ framework [73]. . They performed calculations, especially renormalization, in terms of the coefficients of higher-dimensional operators, and then transformed the results onto the basis of $\kappa_{\lambda_{\rm 3H}}$ and $\kappa_{\lambda_{\rm 4H}}$. These results can not be directly compared with the experimental analysis in the $\kappa$ framework.
@@
Figure 3: Cross sections of Higgs boson pair production at the 13 TeV LHC including both ggF and VBF processes as a function of $\kappa_{\lambda_{\rm 3H}}=\kappa_{\lambda_{\rm 4H}}=\kappa_{\lambda}$. The black line represents the LO result, and the green line denotes the result with (N)NNLO QCD corrections in the ggF (VBF) channel. The red line indicates the result including higher power dependence on the Higgs boson self-coupling. The current and improved upper limits are labeled by the points A and B, respectively.
@@
In Fig. 3, we show different perturbative predictions for the cross sections of both ggF and VBF Higgs boson pair production at the 13 TeV LHC as a function of $\kappa_{\lambda_{\rm 3H}}=\kappa_{\lambda_{\rm 4H}}=\kappa_{\lambda}$. It is evident that higher-order perturbative corrections dramatically change the functional form. The current experimental upper limit by the ATLAS (CMS) collaboration on $\kappa_{\lambda}$ is $6.6$ (6.49) based on the theoretical predictions at QCD NNLO in the ggF channel and NNNLO in the VBF channel; see Table 1. Taking $\delta\sigma^{\kappa_{\lambda}}_{\rm EW}$ corrections into account and assuming that the QCD and EW corrections are factorizable, the upper limit is narrowed down to 5.4 (5.37). These limits are almost the same when keeping $\kappa_{\lambda_{\rm 4H}}=1$. If the scale uncertainties are considered [74], the upper limit spans in the range $(6.5,6.8)$ in the ATLAS result, which would decrease to $(5.4,5.6)$ after including higher power dependence. In the CMS result, the upper limit changes from (6.40, 6.67) to (5.31, 5.48). The lower limits are only modified slightly.
@@
Lastly, the Higgs boson pair invariant mass $m_{HH}$ distributions are shown in Fig. 4. The peak position moves from 400 GeV to 260 GeV when $\kappa_{\lambda}$ varies from 1 to 6, which indicates the Higgs bosons tend to be produced with very small velocity in the case of large $\kappa_{\lambda}$ values. This feature could help to set optimal cuts in the experimental analysis to enhance the sensitivity. From this figure, we also observe that the $\lambda$ dependent corrections have a great impact on the distributions, especially in the small $m_{HH}$ region, and therefore they should be included in future studies.
@@
Figure 4: The Higgs boson pair invariant mass distributions at LO and with $\delta\sigma_{\rm EW}^{\kappa_{\lambda}}$ corrections at the 13 TeV LHC. We have used $\kappa_{\lambda_{\rm 3H}}=\kappa_{\lambda_{\rm 4H}}=\kappa_{\lambda}$.
@@ 4 Conclusion
The precise shape of the Higgs potential stands as a fundamental enigma in particle physics. The current limits on the Higgs self-coupling are predominantly derived based on the assumption that the cross section of the Higgs boson pair production is a quadratic function of the self-coupling. We find that the function form should be generalized to include quartic and cubic power dependencies on the Higgs self-coupling which arise due to higher-order quantum corrections induced by virtual Higgs bosons.
We propose a proper renormalization procedure to explicitly retain the Higgs self-couplings at each calculation step and introduce renormalization of the coupling modifiers to ensure the cancellation of all ultraviolet divergences. We present numerical results of the cross sections of both the ggF and VBF channels at the LHC including higher power dependencies on the Higgs self-coupling. With these improved function forms, we demonstrate that the upper limit set by the ATLAS (CMS) collaboration on the trilinear Higgs self-coupling normalized to its SM value is reduced from 6.6 (6.49) to 5.4 (5.37). This more precise constraint is achieved without more data being analyzed, underscoring the critical importance of incorporating higher power dependencies on the Higgs self-coupling in the cross section.
Furthermore, we find it hard to derive any useful constraint on the quartic Higgs self-coupling solely from Higgs boson pair production. To probe the quartic self-coupling, alternative channels such as triple Higgs boson production may be explored, necessitating collider facilities with higher energies than the LHC to unveil insights into this aspect of the Higgs potential.
@@ Acknowledgments
We express our gratitude to Huan-Yu Bi, Yan-Qing Ma, and Huai-Min Yu for comparing the numerical results of the two-loop amplitudes. We also thank Shan Jin, Yefan Wang and Lei Zhang for helpful discussion. This work was supported in part by the National Natural Science Foundation of China under grant No. 12275156, No. 12321005, No. 12375076 and the Taishan Scholar Foundation of Shandong province (tsqn201909011).